OpenStudy (anonymous):

The difference of a number squared and the number is one less than four times the number.

OpenStudy (anonymous):

n^2-n= 4n -1 n^2 -5n +1 = 0 solve using quadratic equation solution formula

OpenStudy (anonymous):

so its 0?

OpenStudy (anonymous):

no, you need to use the quadratic equation formula ax^2 + bx + c = 0 solved by x = [-b +- sqrt(b^2-4ac)]/2a